OpenStudy (anonymous):

solve using laplace transforms y''+2y'+5y=(e^-t)sint

OpenStudy (irishboy123):

got any IV's?

OpenStudy (anonymous):

use identity for laplace diffrential form

OpenStudy (anonymous):

\(y" +2y' +5y= e^{-t}sin t\) \(\mathcal{L}(y"+2y'+5y) = \mathcal {L}(e^{-t}sin(t)) \) The left hand side : \(\mathcal{L}(y") = s^2\mathcal {L}(y) -sy(0)-y'(0) \\\mathcal{L}(y')=s\mathcal{L}(y)-y(0)\) The \(RHS = \dfrac{1}{(s+1)^2 +1}\). The expression becomes \(s^2\mathcal {L}(y) -sy(0)-y'(0) +2\mathcal{L}(y)-2y(0) + 5\mathcal{L}(y)= \dfrac{1}{(s+1)^2 +1}\) From here, you need y(0) and y'(0) to solve for \(\mathcal {L}(y)\)